OpenStudy (anonymous):

Solve. Please show all of your work. A triangular lot has a perimeter of 1900 feet. One side is 200 feet longer than the shortest side, while the third side is 500 feet longer than the shortest side. Find the lengths of all three sides

OpenStudy (anonymous):

So you have 3 sides of a total of 1900 feet. There is a shortest side, let's call it x because we don't know it yet. One side is 200 feet longer than the shortest, say x + 200. And the last side is 500 feet longer than the shortest, say x + 500. So if you put them together, this should be the equation: x+(x+200)+(x+500)=1900 combining like terms: 3x+700=1900 -700 -700 3x=1200 x=400 That is the length of the shortest side. So the sides are 400, 600, 900.

OpenStudy (anonymous):

i was close lol